OpenStudy (anonymous):

Write the quadratic equation in general form. What is the value of b2 - 4ac? 1 = 2x2 + 5x?

OpenStudy (anonymous):

Do you know the general form for the quadratic?

OpenStudy (anonymous):

do you want the value of b^2 - 4ac or do you want the value of the sqrt(b^2 - 4ac) ?

OpenStudy (anonymous):

Well...if it's zero, the sqrt doesn't matter, so

OpenStudy (anonymous):

the quatdratic formula is b+- squaroot b^2-4ac/2a

OpenStudy (anonymous):

isn't c = -1

OpenStudy (anonymous):

The standard form is: \[ax^2 + bx + c = 0\] So to get the equation you have into that form we should do what?

OpenStudy (anonymous):

25 - 4(2)(-1) = 25 - 8 = 17

OpenStudy (anonymous):

Write the quadratic equation in general form. What is the value of b^2 - 4ac? 1 = 2x^2 + 5x? a17 b33 c34

OpenStudy (anonymous):

thtas the whole question(:

OpenStudy (anonymous):

I think it's a lol

OpenStudy (anonymous):

thank you